# Towards Interactive and Learnable Cooperative Driving Automation: a Large Language Model-Driven Decision-Making Framework

### v2

@@ III-B2 Model-based Control Execution Sub-module
While LLMs possess strong reasoning capabilities, they do not perform well in precise mathematical calculations and low-level vehicle motion control. To address this, we designed a model-based control execution sub-module. In this model, we use a model-based control approach to translate the semantic outputs of the reasoning module into acceleration and front wheel angles, which are then used to update the vehicle’s position, speed, and other information at the next time step. Specifically, both acceleration and front wheel angles are controlled by relatively simple proportional controllers, with the calculation formulas as follows:
@@
where $K_{a}$ and $K_{h}$ are the proportional longitudinal control and control gain, $v_{e}$ and $a_{r}$ are the expect velocity and reference acceleration, $\phi_{r}$ is reference heading to follow the lane, and $\phi$ is the heading.
@@ III-C Reasoning Module
In this subsection, we establish an integrated reasoning module that progresses from state sharing to intent sharing, negotiation, and finally, decision. This module operates in a chain of thought manner, ensuring a smooth transition for CAVs from environmental perception to interaction and negotiation, and ultimately to decision-making.
The reasoning module first extracts information about surrounding vehicles from the environment to create a scene description. It then organizes the states of the vehicles into conflict pairs, forming conflict descriptions. To ensure consistency in vehicle decision-making during conflicts and to avoid collisions, we developed an LLM-based conflict coordinator. This coordinator integrates the current conflict descriptions with traffic rules to determine the order of priority for each conflict group. Finally, each CAV makes its decisions based on the conflict coordinator’s recommendations and its own scenario description.
III-C1 State-perception
The state-perception function is responsible for acquiring and processing information about the current environment of the CAV, including dynamic data such as lane information and vehicle information. The designed state-perception function aligns with the first level of CDA in the SAE J3216 standard, which is state sharing. CAVs are allowed to exchange information with others, therefore paving the way for subsequent higher levels of CDA. Through comprehensive analysis of the above information, the state-perception function can construct a complete and accurate driving environment recognition, which provides a reliable basis for subsequent reasoning.
Specifically, lane information is divided into three categories: the ego lane, adjacent lanes, and conflict lanes. The ego lane is the lane where the ego vehicle is currently driving in, adjacent lanes are the lanes on the left and right sides of the ego lane if they exist, and conflict lanes are the lanes that intersect with the ego lane. Similarly, vehicle information can be grouped based on their relationship with the ego vehicle into leading vehicles, rearing vehicles, surrounding vehicles, and conflict vehicles. During the state-perception process, lane information and vehicle information are combined to create an overview of the ego vehicle’s surrounding environment. Since vehicles in different lanes can affect the ego vehicle differently, a three-level action safety assessment is designed to ensure safe driving, which will be detailed in the decision subsection.
@@ III-C2 Intent-sharing
The intent-sharing function, which conveys the driving intentions of a vehicle to other CAVs, is a key advantage of cooperative driving. From macro to micro, driving intention mainly includes the sharing of expected lanes and expected speed. Through intent-sharing, other vehicles can better understand the ego vehicle’s intentions, allowing them to make decisions while avoiding conflicts as much as possible.
State perception and intent sharing are combined as the scene descriptions. This description is based on the lane and includes all objects that could affect the ego vehicle, such as adjacent vehicles and conflicting vehicles. Therefore, it can be migrated to multiple scenarios, like intersections, highways, and merging areas, significantly enhancing the versatility of the proposed cooperative driving framework.
@@ III-C3 Negotiation
=$=\begin{cases}\Delta TTCP\leq 2s,\quad\text{Serious Danger}\\ 2s<\Delta TTCP\leq 5s,\quad\text{General Danger}\\ 5s<\Delta TTCP\leq 8s,\quad\text{Slight Danger}\\ \Delta TTCP\geq 8s,\quad\text{No Danger}\end{cases}$
where $TTCP$ is the time to conflict point based on the vehicle’s current distance to conflict point $d$ and speed. Among them, when $\Delta TTCP$ is less than 2s, it signifies a serious conflict, requiring at least one of them must take an emergency brake. If $\Delta TTCP$ falls between 2s and 5s, it is considered that there is a general conflict between the two vehicles, and at least one vehicle should slow down and yield. When the $\Delta TTCP$ is between 5s and 8s, the vehicles are considered to have a slight collision, and neither vehicle should accelerate simultaneously. When $\Delta TTCP$ is greater than 8s, it is considered safe that the conflict will not cause a collision.
The conflict coordinator determines the passing sequence of vehicles in each conflict pair based on the severity of the conflict. In this process, traffic rules and social norms in the driving process are simultaneously considered by the conflict coordinator. For example, as shown in Fig. 1, according to traffic rules, turning vehicles should yield to going-straight vehicles. Hence, after negotiation, the conflict coordinator determines that $\text{CAV}_{2}$ should yield in this pair of conflicts. The negotiation result and reasons are sent to the decision function to make the final decision. However, it is important to note that the negotiation result is advisory, and the final decision also depends on other factors in the ego vehicle’s surrounding environment.
@@ III-C4 Decision
The decision module synthesizes the information from the three aforementioned modules to generate the final decision. Considering that the strength of LLMs lies in their reasoning capabilities based on world knowledge rather than numerical computation, we design the decision actions as discrete semantic decisions. The action space $A$ is designed as a set of high-level meta-actions. These meta-actions are then mapped to specific reference acceleration $a_{r}$ and reference heading $\phi_{r}$ values through a numerically based control model. Moreover, the vehicle’s next state is determined based on Eq. 5 and Eq. 6 to ensure safe and stable control.
The decision function mainly includes two tasks: scenario prompt generation and action safety assessment. Scenario prompt generation involves integrating the scene descriptions produced by state perception and intent sharing, as well as the negotiation results from the conflict coordinator, to summarize the key information in the surrounding driving environment. This integration serves as the prompt for LLM decision-making. Furthermore, compared with traditional decision-making methods, the safety of LLM decisions is often difficult to guarantee. Therefore, a three-layer action safety assessment method is proposed to determine whether action would pose safety risks in the current driving environment.
@@ III-D Memory Module
Enhancing continuous learning capabilities in autonomous systems has always been a significant challenge. Novice drivers accumulate experience through continuous driving practice, evaluate the effects of different behaviors, and learn from them to improve their driving skills. Drawing on this mechanism, a memory module is introduced to enable CAVs to learn from past experiences and utilize this knowledge to future interactions. This process is also referred to as Retrieve-Augmented Generation (RAG).
@@ III-D2 Memory retrieval
Specifically, current scenario descriptions and conflict descriptions are converted into vector forms and use cosine similarity to rank memories in the library based on their relevance to the current scenario. The top-ranked memories are then selected as part of the prompts for CAV inference. These similar memories, referred to as few-shot experiences, are injected into the CAV’s reasoning module, enabling the CAV to learn from past mistakes.
@@ IV Experiments and Analysis
To verify whether the proposed CoDrivingLLM can effectively improve the interaction ability and learning ability of CAV, this section carries out verification from three aspects. Firstly, ablation experiments were conducted on the negotiation module in different scenarios. Second, we compared the performance of CoDrivingLLM in different scenarios under 0-shot, 2-shots and 5-shots experiences. Finally, we evaluated the safety and efficiency of CoDrivingLLM against other cooperative driving methods, including optimization-based, rule-based, and machine learning-based methods.
@@ IV-A Experiment Settings
Figure 4: Three scenarios we utilize to train and test our method, (a) scenario 1, highway environment; (b) scenario 2, merge environment; (3) scenario 3, intersection environment.
@@
Simulation Environment. We develop our environment module based on the highway-env[26], an open-sourced and flexible simulation platform for autonomous driving. Three scenarios are designed to conduct experiments, as shown in Fig.4, including a four-way highway scenario, merge scenario, and single-lane unsignalized intersection. The settings of context traffic flow and HDVs follow the instructions of the environment module of our framework.
@@
Implementation details. The GPT-4o mini is utilized our base LLM model to conduct high-level logical thinking and judgment. Four CAVs are controlled in each simulated environment. Furthermore, all scenarios with different settings are repeated 20 times with different random seeds to attain the final results. The success rate is used as the indicator to evaluate the performance of all methods. In one case, if all CAVs finish their driving tasks safely and arrive at their destination, this case is successful. Success rate means the ratio of the number of successful cases to the total number of cases.
IV-B Ablation Study on Negotiation
Lack of interaction ability is an important reason for the current CAV’s struggle on the open road. Cooperative driving, through the sharing of states and intentions, can provide additional information to help conflicting parties align their decision-making. To achieve this, we introduce a negotiation module into the reasoning process. By establishing an LLM-driven conflict coordinator, the traffic rules and scenario description are integrated to produce a suggested traffic sequence to assist the final decision, thus improving the interaction ability. To verify the effectiveness of the conflict coordinator in improving interaction, we conduct ablation experiments comparing the success rate of CoDrivingLLM in various scenarios with and without the negotiation module. In addition, we also select an example case to analyze how the negotiation function can improve safety.
Figure 5: The ablation experiment on negotiation module.
As depicted in Fig.5, in the highway scenario, the presence or absence of the negotiation module does not affect the performance of cooperative driving, as our method consistently achieves a 100% success rate. We attribute this to the lack of significant interaction conflicts in highway scenarios, where the negotiation module does not have an opportunity to intervene effectively. However, in the merge and intersection scenarios, the negotiation module significantly improves the success rate of decision-making. In these two types of scenarios, inevitable conflicts arise between vehicles, and collisions can only be avoided if their decision-making is aligned. Specifically, in the merge scenario, the success rate of CAV decisions increases from 33% to 75%, and in the intersection scenario, it rises from 15% to 78%.
The negotiation module has the most substantial impact in the intersection scenario, where the complexity of interactions and the density of traffic conflicts are much higher than in the other two scenarios. This demonstrates that our negotiation module effectively handles complex traffic conflicts and interaction scenarios, making it a critical component of the overall cooperative driving framework.
@@ IV-B Ablation Study on Negotiation
IV-C Comparison with Memory Module
Figure 7: The performance evaluation of memory module with different shots experiences.
The memory module helps CAVs learn from past experiences to avoid repeating mistakes. To validate the effectiveness of the memory module, we compare the performance of our method in different scenarios under 0-shot, 2-shots, and 5-shots experiences. In this context, 0-shot means the memory module is not used, while 2-shots and 5-shots refer to the introduction of the two and five most similar past experiences to the current scenario as prompts before reasoning.
As shown in Fig.7, in the highway scenario, where vehicle conflicts and interaction challenges are minimal, a 100% safety rate is achieved even without the memory module, and this success rate remains consistent when 2-shots and 5-shots experiences are added. However, in the merge and intersection scenarios, increasing the memory module from 0-shot to 2-shots significantly improves the decision-making success rate, with the merge scenario rising from 78% to 90% and the intersection scenario from 75% to 85%.
We do note that when the memory module is further increased from 2-shots to 5-shots, there is a slight decline in performance in both scenarios. We believe this indicates that longer memory does not necessarily enhance LLM decision-making. An excess of repetitive or low-value memories may occupy too much of the LLM’s context window, reducing its ability to focus on other relevant perceptual information and potentially leading to a degree of negative impact.
@@ IV-C Comparison with Memory Module
We also use an intersection scenario case to demonstrate the role of the memory module. Fig.8(a) shows a typical case of decision failure. In this instance, CoDrivingLLM is not equipped with the memory module. At 4 seconds, although the decision module, through the negotiation module, identifies a potential rear-end conflict between $\text{CAV}_{7}$ and $\text{CAV}_{4}$ and attempts to resolve it by adjusting actions, $\text{CAV}_{4}$ accelerates while $\text{CAV}_{7}$ decelerates. However, due to $\text{CAV}_{7}$’s high initial speed, it collides with $\text{CAV}_{4}$ before it can fully decelerate. This case illustrates that in certain special scenarios, cooperative decision-making can still fail even with the negotiation module in place.
@@
This case effectively showcases the learning utility provided by the memory module to CoDrivingLLM.
@@ IV-D Comparison with Other Methods
Considering that not all cooperative driving methods are suitable for multiple scenario, and unsignalized intersections are often where accidents occur most frequently [27]. Therefore, we selected the three typical cooperative driving methods mentioned earlier to test our approach under unsignalized intersection scenario, including optimization-based method (iDFST)[9], rule-based method (Cooperitave game)[28], learning-based method (MADQN)[29].
@@
Methods Success rate Post Encroachment Time Travel Velocity
@@ IV-D1 Overall Performance
Firstly, we summarized the success rate of each methods. According to the data presented in Table. I, CoDrivingLLM outperforms the other methods, achieving a success rate of 90%. The iDFST and Cooperative Game methods lag slightly behind, both achieving a success rate of 85%. In contrast, MADQN displays the poorest generalization and performance, with a success rate of merely 20%.

### v3

@@ III-B2 Model-based Control Execution Sub-module
While LLMs possess strong reasoning capabilities, they do not perform well in precise mathematical calculations and low-level vehicle motion control. To address this, the target speed and target lane are adjusted based on the semantic output of the reasoning module. Subsequently, a model-based control approach is used to determine the acceleration and front wheel angle, which is formulated as follows:
@@
where $K_{a}$ and $K_{h}$ represent the proportional gains for longitudinal control and lateral control, respectively, with values set to 1.6 and 5, $v_{r}$ and $a_{r}$ are the reference velocity and reference acceleration, $\phi_{r}$ is reference heading to follow the lane, and $\phi$ is the heading.
@@ III-C Reasoning Module
In this subsection, we establish an integrated reasoning module that progresses from state sharing to intent sharing, negotiation, and finally, decision. This module operates in a Chain-of-Thought (CoT) manner, ensuring a smooth transition for CAVs from environmental perception to interaction and negotiation, and ultimately to decision-making.
The reasoning module first extracts information about surrounding vehicles from the environment to create a scene description. It then organizes the states of the vehicles into conflict pairs, forming conflict descriptions. To ensure consistency in vehicle decision-making during conflicts and to avoid collisions, we develop an LLM-based conflict coordinator. This coordinator integrates the current conflict descriptions with traffic rules to determine the order of priority for each conflict group. Finally, each CAV makes its decision based on the conflict coordinator’s recommendations and its own scenario description.
III-C1 State-sharing
The state-perception function is responsible for acquiring and processing information about the current environment of the CAV, including dynamic data such as lane information and vehicle information. The designed state-sharing function aligns with the first level of CDA in the SAE J3216 standard, which is state sharing. CAVs are allowed to exchange information with others, therefore paving the way for subsequent higher levels of CDA. Through comprehensive analysis of the above information, the state-sharing function can construct a complete and accurate driving environment recognition, which provides a reliable basis for subsequent reasoning.
Specifically, lane information is divided into three categories: the ego lane, adjacent lanes, and conflict lanes. The ego lane is the lane where the ego vehicle is currently driving in, adjacent lanes are the lanes on the left and right sides of the ego lane if they exist, and conflict lanes are the lanes that intersect with the ego lane. Similarly, vehicle information can be grouped based on their relationship with the ego vehicle into leading vehicles, rearing vehicles, and conflict vehicles. During the state-sharing process, lane information and vehicle information are combined to create an overview of the ego vehicle’s surrounding environment. Since vehicles in different lanes can affect the ego vehicle differently, a three-level action safety assessment is designed to ensure safe driving, which will be detailed in the decision subsection.
@@ III-C2 Intent-sharing
The intent-sharing function, which conveys the driving intentions of a vehicle to other CAVs, is a key advantage of cooperative driving. From macro to micro, driving intention mainly includes the sharing of expected lane and expected speed. Through intent-sharing, other vehicles can better understand the ego vehicle’s intentions, allowing them to make decisions while avoiding conflicts as much as possible.
State sharing and intent sharing are combined as the scene descriptions. This description is based on the lane and includes all objects that could affect the ego vehicle, such as adjacent vehicles and conflicting vehicles. Therefore, it can be migrated to multiple scenarios, like intersections, highways, and merging areas, significantly enhancing the versatility of the proposed cooperative driving framework.
@@ III-C3 Negotiation
[…]
where $TTCP$ is the time to conflict point based on the vehicle’s current distance to conflict point $d$ and speed $v$. Among them, when $\Delta TTCP$ is less than 2s, it signifies a serious conflict, requiring at least one of them must take an emergency brake. If $\Delta TTCP$ falls between 2s and 5s, it is considered that there is a general conflict between the two vehicles, and at least one vehicle should slow down and yield. When the $\Delta TTCP$ is between 5s and 8s, the vehicles are considered to have a slight collision, and vehicles should not accelerate simultaneously. When $\Delta TTCP$ is greater than 8s, it is considered safe that the conflict will not cause a collision. After discretizing $\Delta TTCP$ into several intervals, the different action risk is described in prompt tailored to the severity of the conflict.
The conflict coordinator determines the passing sequence of vehicles in each conflict pair based on the severity of the conflict. Based on the generated scenario description, information such as the distance to the conflict point and the time to the conflict point is derived to characterize the conflict description. In this process, the conflict coordinator utilizes the conflict description while simultaneously considering traffic rules and social norms in driving to produce the negotiation result. For example, as shown in Fig. 1, according to traffic rules, turning vehicles should yield to going-straight vehicles. Hence, after negotiation, the conflict coordinator determines that $\text{CAV}_{2}$ should yield in this pair of conflicts. The negotiation result and reasons are sent to the decision function to make the final decision. However, it is important to note that the negotiation result is advisory, and the final decision also depends on other factors in the ego vehicle’s surrounding environment.
@@ III-C4 Decision
The decision module synthesizes the information from the three aforementioned modules to generate the final decision. Considering that the strength of LLMs lies in their reasoning capabilities based on world knowledge rather than numerical computation, we design the decision actions as discrete semantic decisions. The action space $A$ is designed as a set of high-level meta-actions. These meta-actions are then mapped to specific reference acceleration $a_{r}$ and reference heading $\phi_{r}$ values through based on control model. Moreover, the vehicle’s next state is determined based on Eq. 2 and Eq. 3 to ensure safe and stable control.
The decision function mainly includes two tasks: scenario prompt generation and action safety assessment. Scenario prompt generation involves integrating the scene descriptions produced by state sharing and intent sharing, as well as the negotiation results from the conflict coordinator, to summarize the key information in the surrounding driving environment. This integration serves as the prompt for LLM decision-making. Furthermore, compared with traditional decision-making methods, the safety of LLM decisions is often difficult to guarantee. Therefore, a three-layer action safety assessment method is proposed to determine whether action would pose safety risks in the current driving environment.
@@ III-D Memory Module
Enhancing continuous learning capabilities in autonomous systems has always been a significant challenge. Novice drivers accumulate experience through continuous driving practice, evaluate the effects of different behaviors, and learn from them to improve their driving skills. Drawing on this mechanism, a memory module is introduced to enable CAVs to learn from past experiences and utilize this knowledge to future interactions. This process is also referred to as RAG.
@@ III-D2 Memory retrieval
Specifically, current scenario descriptions and conflict descriptions are converted into vector forms and cosine similarity is employed to rank memories in the library based on their relevance to the current scenario. The top-ranked memories are then selected as part of the prompts for CAV inference. These similar memories, referred to as few-shot experiences, are injected into the CAV’s reasoning module, enabling the CAV to learn from past mistakes.
@@ IV Experiments and Analysis
To verify whether the proposed CoDrivingLLM can effectively improve the interaction ability and learning ability of CAV, this section carries out verification from three aspects. Firstly, ablation experiments were conducted on the various sub-module in Reasoning Module in different scenarios. Second, the learning ability of CoDrivingLLM was evaluated by comparing its performance across different scenarios after experiencing varying numbers of interactions. Finally, we evaluated the safety and efficiency of CoDrivingLLM against other cooperative driving methods, including optimization-based, rule-based, and machine learning-based methods.
@@ IV-A Experiment Settings
Figure 4: Four scenarios we utilized to train and test our method, (a) scenario 1, intersection environment; (b) scenario 2, roundabout environment; (3) scenario 3, highway environment; (4) scenario 4, merge environment.
@@
Simulation Environment. We develop our environment module based on the highway-env [38], an open-sourced and flexible simulation platform for autonomous driving. Four scenarios are designed to conduct experiments, as shown in Fig.4, including single-lane unsignalized intersection, roundabout scenario, four-way highway scenario, and merge scenario. The settings of context traffic flow and HDVs follow the instructions of the environment module of our framework.
@@
Implementation details. The GPT-4o mini is utilized as our base LLM model to conduct high-level logical thinking and judgment. Furthermore, all scenarios are repeated 20 times using different random seeds to randomly generate the initial positions, speeds, and expected destinations of the vehicles. The success rate is used as the indicator to evaluate the performance of all methods. A case is considered successful if all CAVs finish their tasks safely and reach their destination.
IV-B Ablation Study on Sub-modules of the Reasoning Module
Cooperative driving has the potential to enhance the performance of CAVs in complex scenarios. Following the SAE J3216 standard, we developed a reasoning module that integrates state sharing, intention sharing, negotiation, and decision. To validate the effectiveness of these sub-modules, ablation studies were conducted in this subsection.
When the state-sharing sub-module is disabled, CAVs can only access the position and speed information of nearby vehicles within a 50-meter range. Moreover, CAVs are unable to detect potential conflicts between other vehicles and themselves without the intention-sharing sub-module. Finally, when the negotiation sub-module is absent, each CAV independently generates its expected passing order based on its understanding of the scenario information, rather than relying on a centralized conflict coordinator.
Figure 5: The ablation experiment on sub-modules of the reasoning module.
As shown in Fig.5, the negotiation module has the most significant impact on model performance. In roundabout and intersection, removing the negotiation module reduces the success rate to less than 15%, which is less than one-fifth of the original model’s performance. This result underscores the negotiation module’s effectiveness in resolving complex traffic conflicts, making it a vital component of the cooperative driving framework.
Additionally, the state-sharing module has a comparatively smaller effect. In all scenarios, the success rate remains above 50%. It performs especially well in merging zones, where it closely matches the performance of CoDrivingLLM. This is due to the fact that vehicles in merging scenarios often travel parallel for extended stretches, allowing for earlier detection of interacting objects in adjacent lanes and more timely responses.
Finally, the intention-sharing module’s impact varies across scenarios. In environments with concentrated conflict points, such as merging zones and intersections, its influence is relatively minor. However, in roundabouts, where conflict points are more dispersed and frequent, the absence of state sharing leads to a significant decline in success rates. Meanwhile, in highway scenarios, where car-following interactions dominate over conflicts, the success rate remains consistently high regardless of the module’s presence.
@@ IV-B Ablation Study on Sub-modules of the Reasoning Module
IV-C The Performance of Continuous Learning
The memory module enables CAVs to learn from past experiences, preventing the repetition of mistakes. To validate the continuous learning capability of the proposed framework, we compared the changes in the model’s success rate after different numbers of interactions.
Figure 7: The performance evaluation of memory module with different interaction numbers.
In the highway environment, as shown in Fig.7, the success rate reaches 100% after just 5 interactions. This high success rate can likely be attributed to the relatively simple conflict patterns and lower complexity compared to other scenarios. In the merging scenario, the success rate reaches around 75% after 10 interactions. However, intersections and roundabouts show lower success rates during the initial interactions due to their higher complexity, which includes a variety of conflict types and multiple conflict points. Overall, the success rate improves significantly with the number of interactions across various scenarios, demonstrating that the memory module enables CoDrivingLLM to effectively learn from interactions.
@@ IV-C The Performance of Continuous Learning
We also used an intersection scenario case to demonstrate the role of the memory module. Fig.8(a) shows a typical case of decision failure. In this instance, CoDrivingLLM is not equipped with the memory module. At 4 seconds, although the decision module, through the negotiation module, identifies a potential rear-end conflict between $\text{CAV}_{7}$ and $\text{CAV}_{4}$ and attempts to resolve it by adjusting actions, $\text{CAV}_{4}$ accelerates while $\text{CAV}_{7}$ decelerates. However, due to $\text{CAV}_{7}$’s high initial speed, it collides with $\text{CAV}_{4}$ before it can fully decelerate. This case illustrates that in certain special scenarios, cooperative decision-making can still fail even with the negotiation module in place.
@@ IV-D Comparison with Other Methods
Considering that not all cooperative driving methods are suitable for multiple scenarios, and unsignalized intersections are where accidents occur most frequently [39]. Therefore, we selected the six typical cooperative driving methods mentioned earlier to test our approach under unsignalized intersection scenario, including optimization-based method (Cooperative game [40], MCTS [41]), rule-based method (iDFST [13], FCFS), learning-based method (MADQN [42]), and LLM-based (Dilu [33]).
@@
[…]
@@ IV-D1 Overall Performance
Firstly, we evaluated the overall performance of different methods based on their success rates, as summarized in Table. I. The proposed CoDrivingLLM and the rule-based method FCFS both achieve a 90% success rate across cases with various initial states, the highest among all methods. However, our method demonstrates a significant advantage in efficiency compared to FCFS, which will be discussed in detail below.
In addition, the rule-based method iDFST achieves a success rate of 85%. Among optimization-based approaches, MCTS and Cooperative game achieves success rates of 65% and 55%, respectively, placing them in the mid-range of performance. Moreover, although Dilu is a representative work that employs LLM for decision-making in CAV, it was originally designed for highway environments and not for cooperative driving. As a result, when applied to intersection scenarios, its success rate drop to 38%. Lastly, MADQN displays the poorest generalization and performance, with a success rate of only 20%.